I have one more question. Could you please show the steps?

A positive integer is 5 less than another. If the reciprocal of the smaller integer is subtracted from twice the reciprocal of the larger, then the result is 7/204. Find the two integers.

Answer by mananth(16946)

A positive integer is 5 less than another.
let one integer be x
the other (x-5)
the reciprocal of the smaller integer is subtracted from twice the reciprocal of the larger,
2/x - 1/(x-5) = 7/204
(2(x-5)-x)/x(x-5) =7/204
(2x-10-x)/(x^2-5x) =7/204
(x-10)*204 = 7(x^2-5x)
204x -2040 =7x^2 -35x
7x^2-239x +2040=0
Solve the quadratic
the numbers are 17 & 12